# The equation of Binet in classical and relativistic orbital mechanics

### v1

@@ Abstract
Binet’s equation provides a direct way to obtain the geometric shape of orbits in a central force field. It is well known that in Newtonian gravitation Binet’s equation leads to all the conic curves as solutions for an inverse-square force. In this work, we show how Binet’s equation arises from the horizontal and vertical infinitesimal displacements of a body in free fall and in inertial motion. This derivation uses elementary concepts of infinitesimal calculus. Second, we derive the relativistic version of Binet’s equation for the Schwarzschild-(anti-)de Sitter metric. This derivation, which is novel, directly relates the coordinates involved in Binet’s equation without the need to introduce potentials or the use of Killing vectors. Finally, we tackle some controversies related to the role of the cosmological constant in the trajectory of photons in a Schwarzschild-(anti-)de Sitter or even in Reissner–Nordström–(anti-)de Sitter spacetimes.
@@ I Introduction
Among many others, two of Newton’s foremost contributions were the unification of the phenomena of the fall of heavy objects and the motion of celestial bodies, on the one hand, and the systematization of earlier infinitesimal methods into the foundations of modern calculus, on the other. In regard to the motion of celestial bodies, at the very start of his Principia [newton:1687], Newton described the motion of the Moon as a superposition of the motions due to inertia and gravitational fall. As for the infinitesimal methods also used in the Principia, he used them under the idea of the so-called mathematical ultimate equality [needham:2021]. The modern use of these two leads naturally to the equations of motion in the two-body problem, as presented here.
@@
The classical study of orbital motion in a two-body context under central forces is elegantly done with Binet’s equation [gonzalez-lopez:2025]. This differential equation directly yields the geometric shape of the orbit without using time as a parametrization of such two-dimensional curves. It is named after Jacques Philippe Marie Binet (1786-1856), though it is often left unnamed in classical textbooks [symon:71, p. 130][goldstein:2001, p. 87]. For an inverse-square force, all conic curves are solutions of Binet’s equation. The Newtonian Binet’s equation has a counterpart in general relativity which includes additional terms. Several approaches exist to derive it in either context, Newtonian [symon:71, goldstein:2001] or Einsteinian [misner:73, carroll:2019, schutz:2022]. Some original approaches have also been developed, such as in [deliseo:2007], which employs a complex-variable formalism. In the present article we first obtain the non-relativistic Binet’s equation by making use of the two anticipated elements: inertia and the central force producing the dynamics, Newton’s gravitational force in particular. The coordinate system is chosen to underline the idea of a vertical fall onto the attracting body and a given inertial speed. Heretofore, the usual approach is to obtain it from the one-body equivalent problem, written from the start in polar coordinates. Second, we derive the relativistic Binet’s equation by directly relating the coordinates involved in it, that do not include time. This is different from the usual procedures and sheds light both on the tools of general relativity and its flexibility to adopt different coordinate systems.
@@ II Binet’s equation as a consequence of a vertical fall
Figure 1: Representation of the instantaneous position of the reduced mass $\mu$ in a Cartesian coordinate system $(x,y)$ centred at the attracting centre. The vertical $y$ axis is aligned with the direction from $\mu$ to the attracting centre at this instant. The horizontal $x$ axis is perpendicular to $y$. The velocity components $v_{x}$ and $v_{y}$ represent the inertial motion along $x$ and the vertical fall along $y$, respectively. The dotted curves represent possible orbits of $\mu$ around the attracting centre, which will be one or the other depending on the initial conditions.
@@
which is Binet’s equation. Once solved for $u$ and then converted to $r$, it has all the conic curves as solutions:
@@
where $v_{x}=\dot{x}=r\,\dot{\phi}$, $v_{y}=\dot{y}=\dot{r}$. The values of $y$, $v_{x}$ and $v_{y}$ are, by definition, initial values because the Cartesian coordinate system is considered and used here only at $t\in[0,dt]$. It is straightforward from (14) and (15) to see that $y$, $v_{x}$ and $v_{y}$ determine $e$ and $\phi_{0}$  [^3] Alternatively, the integration constants can be obtained from the angular momentum and the energy since $J=y\,v_{x}$ and $E=\frac{1}{2}(v_{x}^{2}+v_{y}^{2})-GM/y$.
Mathematically, the most interesting features of the Binet’s equation for the Newton attractive force are the following:
@@
    It is a linear ordinary differential equation (ODE), unlike the equation in $r=r(t)$ resulting from (9) after introducing (7) and the Newtonian gravitational force
@@
    The term $J^{2}/r^{3}$ encodes the centrifugal force term in (16). The presence of a fictitious force in (16) is due to the fact that we employ a rotating reference frame. The only axis in use is that of the $r$ coordinate, which follows the movement of the reduced mass along the orbit.
@@
    The problem has been reduced to a one-dimensional problem with respect to the azimuth angle, specifically the one of a free, undamped harmonic oscillator. The domain of $u=1/r$ is the positive real semi-axis, which makes it a half-line harmonic oscillator. If $0\leq e<1$, there is a periodic oscillation around $r=J^{2}/GM$, whereas for $e\geq 1$ there is a single pass trajectory. In this case, $u$ approaches 0, which corresponds to $r$ going to infinity. Any initial condition having $u(t_{0})>0$ always produces a positive solution for $r(\phi)$ because, to become negative, $u$ should traverse 0 ($r=\infty$).
@@
    If the initial conditions were $\tilde{u}_{0}=0$ and $\dot{\tilde{u}}_{0}=0$, the trajectory of the particle would be circular. All inertial information has become hidden in the change of variables. For non-circular orbits, the particle’s position oscillates around $\tilde{u}_{0}=0$ in an ellipse or moves about it in a single pass in the cases of a parabola or a hyperbola.
In the next section, we will derive Binet’s equation for the Schwarzschild-(anti-)de Sitter metric. Instead of a force-driven description, the motion is governed by the geodesic equations. A massive body follows a timelike geodesic whose four-velocity has norm $c$, whereas a photon follows a null geodesic whose tangent vector has zero norm. The geometry of the trajectory will not follow from the described combination of a horizontal momentum plus a vertical push. Instead, it will obey a geodesic path relating the first and second order derivatives of the position through the geometry of spacetime. This geometry is determined by a metric whose form is obtained as a solution of Einstein’s field equations.
III Binet’s equation with Schwartzschild-(anti) de Sitter metric
As is well known, the physical meaning of the motion of an object in a gravitational field changes fundamentally in general relativity compared to Newton’s theory. Two of the most important differences are the shift from the concept of force to that of spacetime curvature, and the nonlinearity implied by Einstein’s field equations. The latter means that, unlike in Newtonian gravitation, the two-body problem does not reduce to a one-body problem because the superposition principle does not hold. However, if one of the masses is much larger than the other, it is possible to treat it as a one-body problem where the smaller mass moves in the gravitational field created by the larger, stationary mass. This is the scenario we consider here. Regarding the first point, the concepts of force and inertia are replaced by the motion of the object along a geodesic in curved spacetime. Such a geodesic is determined by the metric that solves Einstein’s equations for the gravitational field created by the massive body.
@@ III Binet’s equation with Schwartzschild-(anti) de Sitter metric
where $R_{\mu\nu}$ is the Ricci curvature tensor, $R$ the Ricci scalar, $g_{\mu\nu}$ the metric tensor, $T_{\mu\nu}$ the stress-energy tensor, $G$ the gravitational constant, and $c$ the speed of light in vacuum  [^7] We use physical units in this article, even though it is more common to choose geometrized units in which $G=c=1$ so that they do not appear in the equations. The latter option yields masses measured in length units. To recover physical units one replaces $M$ by $GM/c^{2}$. In the vacuum outside the massive body, $T_{\mu\nu}=0$ and the solution is given by the Schwarzschild-(anti-)de Sitter metric, namely  [^8] The Schwarzschild–(anti)–de Sitter (often called Kottler) solution is the most general static, spherically symmetric vacuum solution of Einstein’s equations with a nonzero cosmological constant
@@
where $r_{S}$ is the Schwarzschild radius. The function $f(u)$ is not simply an abbreviation for the coefficient of $c^{2}dt^{2}$ in the metric. Noting that $d\tau^{2}=-1/c^{2}\,ds^{2}$ defines the proper time, we see that $f(u)$ represents the rate at which a stationary clock ticks as measured by a distant observer, $d\tau=\sqrt{f(u)}\,dt$  [^9] This is the well-known effect of a clock running slower as it approaches an attracting mass.. After computing the Christoffel symbols for this metric (see Appendix C), the non-affine geodesic equations take the form:
@@
The procedure is now to use (29a) and integrate it as explained in Appendix D to produce (53), which allows $dt/d\phi$ to be replaced in (29b). In addition to this, the square of $du/d\phi$ in the latter can be calculated from the pseudo-normalization condition (45), with $\alpha=\phi$ and $\lambda=\tau$, and again from (53) as in Appendix D, specifically equation (54). The full substitution of these terms yields  [^10] The involved expressions in Appendices D and E contained an integration constant $B$ that is not in (30) as it gets canceled along the way.
@@
If we now substitute $f(u)$ into the Schwarzschild-(anti-)de Sitter metric, we obtain
@@
These relativistic Binet’s equations are usually derived in textbooks by differentiating equation (56) with respect to $\phi$. In fact, as explained in Appendix F, it is (56) that is known to be solved analytically and not (31). This seems to put Binet’s equation behind (54) in importance. However, it is still relevant to have it at hand. First, it is interesting to confront the relativistic version of Binet’s equation with the non-relativistic one. Thus, we note the presence of the relativistic correction terms proportional to $u^{2}$ and $u^{-3}$. These terms are responsible for the non-closure of elliptical orbits and the consequent precession of the perihelion in an elliptical orbit in general relativity. This can be seen qualitatively from the point of view of Bertrand’s theorem. This theorem states that the only central forces in the general Binet’s equation given in (1) that yield closed and bound orbits, such as circles or ellipses, are Newton’s inverse-square law and the Hooke’s law forces. To obtain a quantitative estimate of this effect, we could consider the case of low-eccentricity elliptical orbits. It is straightforward to show that a linearization of the nonlinear terms leads to a harmonic oscillator whose angular period differs from $2\pi$, implying that the orbit is not closed. In effect, by assuming that the ellipse is a deformation of a circle of inverse radius $u_{0}$, we can write
@@
The deviation of $A$ from unity gives rise to the precession of the perihelion.
@@
where $\wp(z)$ is the Weierstrass elliptic function and $\phi_{in}$ is an azimuthal offset for this function. This offset is such that the initial conditions, $(r_{0},\phi_{0})$ are satisfied, i.e. such that $r(\phi_{0})=r_{0}$ (cf. equations (13), (14) and (15)). The solution for the massive particle is significantly more involved and can be found in [hackmann:2008, p. 7, eq. (63)]. Regardless of whether (56) is used for computing it, it is also a solution of (31) for the lightlike case, as it should be. As for the aforementioned controversy, equation (34) depends on $\Lambda$ through the initial conditions. There are other hidden dependencies, such as the one upon the integration constant $B$ and therefore upon the energy and the angular momentum. It is also instructive in this regard to examine (17) in the non-relativistic case and observe the absence of the parameters $GM$ and $J$ in it. Nevertheless, the solution of (17) depends on these parameters through the initial conditions, which determine the eccentricity $e$ and the semi-major axis $a$ of the conic section.
@@ IV Conclusions
In Newtonian mechanics, inertial frames are usually preferred because Newton’s laws hold in them without the introduction of fictitious forces. In general relativity, by contrast, affine parametrizations are recommended for rather different reasons. Both choices share the aim of avoiding additional and somewhat non-physical terms in the equations of motion. Nonetheless, we have used non-inertial frames in Newtonian mechanics and non-affine parametrizations in general relativity to derive Binet’s equation. This procedure is advantageous in the case of the non-relativistic Binet’s equation because it reproduces the historical derivation and the original physical intuition behind it. It begins from a differential analysis of the trajectory in Cartesian coordinates and converts it to polar coordinates, which constitute the most convenient system for central-force problems. In the relativistic case, the use of the inverse radius and the azimuthal coordinate as a non-affine parameter from the outset allows for a direct derivation of Binet’s equation in terms of the coordinates that were found to be ideal in the Newtonian case.
Binet’s equation in its most general form is a one-degree-of-freedom autonomous conservative oscillator  [^12] An oscillator described by $d^{2}\xi/d\alpha^{2}=\sum_{n=-M}^{N}\xi^{n}\;;M,N\in\mathbb{N}$ is autonomous for not including the dependent variable $\alpha$ and conservative because of the integrability of the Laurent polynomial., both in the non-relativistic and relativistic cases. This property makes it particularly suitable for various types of analysis of the motion, such as those based on phase-space diagrams, stability analysis of circular orbits, and perturbation methods for nearly circular orbits. These techniques are widely used in celestial mechanics and dynamical systems theory to understand the behavior of orbits under various conditions. Nevertheless, we have offered a complementary perspective wherefrom conservation quantities such as the energy and angular momentum arise naturally from the geodesic equations with no need to invoke potentials or Killing vectors.

### v2

@@ Abstract
Binet’s equation provides a direct way to obtain the geometric shape of orbits in a central force field. It is well known that in Newtonian gravitation Binet’s equation leads to all the conic curves as solutions for an inverse-square force. Its standard derivation in higher-education textbooks relies on the use of polar coordinates and the conservation of angular momentum and energy. Here an alternative derivation is presented that emphasizes the physical interpretation of the problem by considering the motion as a superposition of a vertical fall toward the attracting centre plus an inertial motion in the horizontal and vertical directions. That is, we show how Binet’s equation arises from the horizontal and vertical infinitesimal displacements of a body in free fall and in inertial motion. This deduction uses elementary concepts of infinitesimal calculus and connects closely with the original geometric and kinematical viewpoint of Newton and his predecessors and contemporaries. It is a view that is fully consistent with the maths taught in undergraduate courses. Second, we derive the relativistic version of Binet’s equation for the Schwarzschild-(anti-)de Sitter metric. This derivation, which is novel, directly relates the coordinates involved in Binet’s equation without the need to introduce potentials or the use of Killing vectors. It is also an adequate step forward in the use of mathematical physics that can be introduced in advanced undergraduate or graduate courses of general relativity. Finally, we tackle some controversies related to the role of the cosmological constant in the trajectory of photons in a Schwarzschild-(anti-)de Sitter or even in Reissner–Nordström–(anti-)de Sitter spacetimes. In this last step of the work, it is shown to students how a thorough yet concise mathematical analysis can solve disputes that arise even in high-level research articles. At the same time, the attentive reader is invited to connect the arguments provided at this final point with the analysis made in the case of non-relativistic Binet’s equation.
@@ I Introduction
Among many others, two of Newton’s foremost contributions were the unification of the phenomena of the fall of heavy objects and the motion of celestial bodies, on the one hand, and the systematization of earlier infinitesimal methods into the foundations of modern calculus, on the other. In regard to the motion of celestial bodies, at the very start of his Principia [newton:1687], Newton described the motion of the Moon as a superposition of the motions due to inertia and gravitational fall. As for the infinitesimal methods used in the Principia, he used them under the idea of the so-called mathematical ultimate equality [needham:2021]. The modern use of these two leads naturally to the equations of motion in the two-body problem, as presented here.
@@
The classical study of orbital motion in a two-body context under central forces is elegantly done with Binet’s equation [gonzalez-lopez:2025]. This differential equation directly yields the geometric shape of the orbit without using time as a parametrization of the two-dimensional curves. It is named after Jacques Philippe Marie Binet (1786-1856), though it is often left unnamed in classical textbooks [symon:71, p. 130][goldstein:2001, p. 87]. For an inverse-square force, all conic curves are solutions of Binet’s equation. The Newtonian Binet’s equation has a counterpart in general relativity which includes additional terms. Several approaches exist to derive it in either context, Newtonian [symon:71, goldstein:2001] or Einsteinian [misner:73, carroll:2019, schutz:2022]. Some original approaches have also been developed, such as in [deliseo:2007], which employs a complex-variable formalism. In the present article we obtain first the non-relativistic Binet’s equation by making use of the two anticipated elements: inertia and the central force producing the dynamics, Newton’s gravitational force in particular. The coordinate system is chosen to underline the idea of a vertical fall onto the attracting body and a given inertial speed. Traditionally, the problem is formulated from the outset in polar coordinates. Here, however, we begin by using a Cartesian coordinate system. Second, we derive the relativistic Binet’s equation by directly relating the coordinates involved in it, that do not include time. This is different from the usual procedures and sheds light both on the tools of general relativity and its flexibility to adopt different coordinate systems.
@@ II Binet’s equation as a consequence of a vertical fall
Figure 1: Representation of the instantaneous position of the reduced mass $\mu$ in a Cartesian coordinate system $(x,y)$ centred at the attracting centre. The vertical $y$ axis is aligned with the direction from the attracting centre to $\mu$ at this instant. The horizontal $x$ axis is perpendicular to $y$. The initial velocity components $v_{x}$ and $v_{y}$ represent the inertial motion along $x$ and $y$, respectively. The dotted curves represent some possible orbits of $\mu$ around the attracting centre, depending on the initial conditions.
@@
which is the Newtonian Binet’s equation. Once solved for $u$ and then converted to $r$, it has all the conic curves as solutions:
@@
where $v_{x}=\dot{x}=r\,\dot{\phi}$, $v_{y}=\dot{y}=\dot{r}$. The values of $y$, $v_{x}$ and $v_{y}$ are, by definition, initial values because the Cartesian coordinate system is considered and used here only at $t\in[0,dt]$. It is straightforward from (14) and (15) to see that $y$, $v_{x}$ and $v_{y}$ determine $p$, $e$ and $\phi_{0}$  [^3] Alternatively, the integration constants can be obtained from the angular momentum and the energy since $J=y\,v_{x}$ and $E=\frac{1}{2}(v_{x}^{2}+v_{y}^{2})-GM/y$. The value of $GM$ determines the numerical mapping of the initial conditions onto these three parameters.
Mathematically, the most interesting features of the Binet’s equation for the Newtonian attractive force are the following:
@@
    It is a linear ordinary differential equation (ODE), unlike the equation in $r=r(t)$ resulting from (9) when introducing (7) and Newton’s gravitational force,
@@
    The term $J^{2}/r^{3}$ encodes the centrifugal force term in (16). The presence of a fictitious force in (16) is due to the fact that we employ a rotating reference frame  [^4] One often finds recommendations on how to deal with fictitious forces that lack full rigour but exhibit a certain degree of practicality. For example, the text on Analytical Mechanics by Hand and Finch [handfinch:98] recommends the following approach: “Treat fictitious forces like real forces, and pretend that you are in an inertial frame”. Such prescriptions, however, are not conceptually optimal.
@@
    The problem has been reduced to a one-dimensional problem with respect to the azimuthal angle, specifically the one of a free, undamped harmonic oscillator. The domain of $u=1/r$ is the positive real semi-axis, which makes it a half-line harmonic oscillator. If $0\leq e<1$, there is a periodic oscillation around $r=J^{2}/GM$, whereas for $e\geq 1$ there is a single pass trajectory. In the latter case, $u$ approaches 0, which corresponds to $r$ going to infinity. Similarly, any initial condition having $u(t_{0})>0$ always produces a positive solution for $r(\phi)$ because, to become negative, $u$ should traverse 0 ($r=\infty$).
@@
    If the initial conditions were $\tilde{u}_{0}=0$ and $\dot{\tilde{u}}_{0}=0$, the trajectory of the particle would be circular. Both $J$ and $GM$ have been absorbed into the change of variables. For non-circular orbits, the particle’s position oscillates around $\tilde{u}_{0}=0$ in an ellipse or moves about it in a single pass in the cases of a parabola or a hyperbola.
In the next section, we will derive Binet’s equation for the Schwarzschild-(anti-)de Sitter metric. Instead of a force-driven description, the motion is governed by the geodesic equations. A massive body follows a timelike geodesic whose four-velocity has norm $c$, whereas a photon follows a null geodesic whose tangent vector has zero norm. The geometry of the trajectory will not follow from the described combination of a horizontal and vertical momentum plus a vertical push. Instead, it will obey a geodesic path relating the first and second order derivatives of the position through the geometry of spacetime. This geometry is determined by a metric whose form is obtained as a solution of Einstein’s field equations.
III Binet’s equation with Schwarzschild-(anti) de Sitter metric
As is well known, the physical meaning of the motion of an object in a gravitational field changes fundamentally in general relativity compared to Newton’s theory. Two of the most important differences are the shift from the concept of force to that of spacetime curvature, and the nonlinearity implied by Einstein’s field equations. The latter means that, unlike in Newtonian gravitation, the two-body problem does not reduce to a one-body problem because the necessary superposition principle does not hold. However, if one of the masses is much larger than the other, it is possible to treat it as a one-body problem where the smaller mass moves in the gravitational field created by the larger, stationary mass. This is the scenario we consider here. Regarding the first point, the concepts of force and inertia are replaced by the motion of the object along a geodesic in curved spacetime. Such a geodesic is determined by the metric that solves Einstein’s equations for the gravitational field created by the massive body.
@@ III Binet’s equation with Schwarzschild-(anti) de Sitter metric
where $R_{\mu\nu}$ is the Ricci curvature tensor, $R$ the Ricci scalar, $g_{\mu\nu}$ the metric tensor, $T_{\mu\nu}$ the stress-energy tensor, $G$ the gravitational constant, and $c$ the speed of light in vacuum  [^8] We use physical units in this article, even though it is more common to choose geometrized units in which $G=c=1$ so that they do not appear in the equations. The latter option yields masses measured in length units. To recover physical units one replaces $M$ by $GM/c^{2}$. In the vacuum outside the massive body, we have $T_{\mu\nu}=0$ and the solution is given by the Schwarzschild-(anti-)de Sitter metric, namely  [^9] The Schwarzschild–(anti)–de Sitter solution is also called Kottler solution.
@@
where $r_{S}$ is the Schwarzschild radius. The function $f(u)$ is not simply an abbreviation for the coefficient of $c^{2}dt^{2}$ in the metric. Noting that $d\tau^{2}=-1/c^{2}\,ds^{2}$ defines the proper time, we see that $f(u)$ determines the relation between this proper time and the coordinate time increment $dt$ at a fixed spatial point, $d\tau=\sqrt{f(u)}\,dt$, i.e. when $du$, $d\theta$, and $d\phi$ are all zero  [^10] This is the well-known effect that a clock runs more slowly at smaller radial distances from a gravitational mass.. After computing the Christoffel symbols for this metric (see Appendix C), the non-affine geodesic equations take the form:
@@
The procedure is now to use (29a) and integrate it as explained in Appendix D to produce (53), which allows $dt/d\phi$ to be replaced in (29b). In addition to this, the square of $du/d\phi$ in the latter can be calculated from the pseudo-normalization condition (45), with $\alpha=\phi$ and $\lambda=\tau$ (or simply $\lambda$ for photons), and again from (53) as in Appendix D, producing equation (54). After all these substitutions, equation (29b) becomes  [^11] The involved expressions in Appendices D and E contained an integration constant $B$ that is not in (30) as it gets canceled along the way.
@@
If we now substitute $f(u)$ into the Schwarzschild-(anti-)de Sitter metric, as well as inserting (52121212The only input from the affine geodesic equations comes from the conservation law given by (52). It effectively compensates for the disappearance of equation (28c)., we obtain
@@
These relativistic Binet’s equations are usually derived in textbooks by differentiating equation (56) with respect to $\phi$. In fact, as explained in Appendix F, equation (56) is of a type for which a general analytical solution is known, unlike equation (31). This seems to put Binet’s equation behind (54) in importance. However, it is still relevant to have it at hand. First, it is interesting to confront the relativistic version of Binet’s equation with the non-relativistic one. Thus, we note the presence of the relativistic correction terms proportional to $u^{2}$ and $u^{-3}$. These terms are responsible for the non-closure of elliptical orbits and the consequent precession of the perihelion in an elliptical orbit in general relativity. This is consistent with Bertrand’s theorem [goldstein:2001]. This theorem states that Newton’s inverse-square and Hooke’s laws are the only central forces in equation (1) for which all bounded orbits are closed. To obtain a quantitative estimate of this effect, we could consider the case of low-eccentricity elliptical orbits. It is straightforward to show that a linearization of the nonlinear terms leads to a harmonic oscillator whose angular period differs from $2\pi$, implying that the orbit is not closed. In effect, by assuming that the ellipse is a deformation of a circle of inverse radius $u_{0}$, we can write
@@
As anticipated, the deviation of $A$ from unity gives rise to the precession of the perihelion.
@@
where $\wp(z)$ is the Weierstrass elliptic function and $\phi_{in}$ is an azimuthal offset for this function. This offset is such that the initial conditions are satisfied (in a similar way to equations (13), (14) and (15) for essentially the same mathematical problem). The solution for the massive particle is significantly more involved and can be found in [hackmann:2008, p. 7, eq. (63)]. As for the aforementioned controversy, equation (34), which is the solution of both (31) and (54) in the lightlike case, depends on $\Lambda$ (see Appendix F). Equation (31) has other hidden dependencies apart from $\Lambda$, such as the one upon the integration constant $B$ and therefore upon the energy and the angular momentum. The initial conditions map onto the solutions of the equations of motion, or geodesic equations in this case, through these apparently hidden constants such as $\Lambda$. It is also instructive in this regard to examine (17) in the non-relativistic case and observe the absence of the parameters $GM$ and $J$ in it. Nevertheless, the solution of (17) depends on them through the definition of the measured quantity, $\tilde{u}$, and through the manner in which the initial conditions determine the eccentricity, $e$, the semilatus rectum, $p$, and the angle of the periapsis, $\phi_{0}$.
@@ IV Conclusions
In Newtonian mechanics, inertial frames are usually preferred because Newton’s laws hold in them without the introduction of fictitious forces. In general relativity, affine parametrizations are recommended for rather different reasons. However, both choices share the aim of avoiding additional and somewhat non-physical terms in the equations of motion. Nonetheless, we have used non-inertial frames in Newtonian mechanics and non-affine parametrizations in general relativity to derive Binet’s equation. This procedure is advantageous in the case of the non-relativistic Binet’s equation because it reproduces the historical derivation and the original physical intuition behind it. It begins from a differential analysis of the trajectory in Cartesian coordinates and converts it to polar coordinates, which constitute the most convenient system for central-force problems. In the relativistic case, the use of the inverse radius and the azimuthal coordinate as a non-affine parameter from the outset allows for a direct derivation of Binet’s equation in terms of the coordinates that were found to be ideal in the Newtonian case.
Binet’s equation in its most general form is a one-degree-of-freedom autonomous conservative ODE  [^14] An ODE described by $d^{2}\xi/d\alpha^{2}=\sum_{n=-M}^{N}\xi^{n}\;;M,N\in\mathbb{N}$ is autonomous for not including the dependent variable $\alpha$ and conservative because of the integrability of the Laurent polynomial., both in the non-relativistic and relativistic cases. This property makes it particularly suitable for various types of analysis of the motion. We have offered a complementary perspective for the relativistic case in which conservation quantities such as the energy and angular momentum arise naturally from the geodesic equations with no need to invoke potentials or Killing vectors.